OpenStudy (vera_ewing):

Which polynomial identity will prove that 72 = 64 + 8?

OpenStudy (vera_ewing):

@Data_LG2

OpenStudy (anonymous):

I don't get it, is there any choices or any other information besides that?

OpenStudy (jtvatsim):

Perhaps they are getting at x(x+1) = x^2 + x

OpenStudy (jtvatsim):

See how the 64 is just 8^2 and 8 is sitting right next to the 64.

OpenStudy (vera_ewing):

Difference of Squares Difference of Cubes Sum of Cubes Square of a Binomial

OpenStudy (vera_ewing):

Maybe it's C?

OpenStudy (anonymous):

I think so.. 64 = 4^3 and 8=2^3 4^3 + 2^3= 72